# Aircraft Trajectory Dataset Augmentation in Latent Space

### v1

@@ 4.2 Model Configurations
We set the number of encoder layers to $L=3$, the hidden dimension to $D=128$, the number of heads to $h=4$, the feedforward dimension to $D_{ff}=512$, and the dropout probability to $p=0.2$. For the training process, we used the ADAM optimizer (Kingma & Ba, 2017) with an initial learning rate of $10^{-4}$. The number of principal components $P$ for PCA was 22, which accounts for 99% of the variance in the dataset. The number of Gaussian distributions $K$ of the GMM was determined as 32 based on the Bayesian Information Criterion (Schwarz, 1978).
@@ 4.3 Comparison Baselines
In this work, we compared the proposed framework with several competitive baselines: TimeGAN (Yoon et al., 2019), TimeVAE (Desai et al., 2021), SMOTE-Interpolation (SMOTE-I), SMOTE-Extrapolation (SMOTE-E) (DeVries & Taylor, 2017), GMM (Barratt et al., 2018) and TCN-VAE (Krauth et al., 2023). We utilized the following publicly available source code for the implementation of the baselines (TimeGAN, TimeVAE, TCN-VAE and GMM). Upon the publication of this paper, our source code for ATRADA will also be made publicly available.
For TimeGAN, the hidden dimensions of its three GRU layers were set to four times the size of the input features. This setting is consistent with the implementation described in the original TimeGAN paper (Yoon et al., 2019). Moreover, the hidden dimensions of TimeVAE and TCN-VAE were configured to match the hidden dimension used in ATRADA. For both SMOTE-I and SMOTE-E (DeVries & Taylor, 2017), we set the number of nearest neighbors to 10, with the interpolation and extrapolation degrees set at 0.5. Given that SMOTE-I and SMOTE-E were tailored to generate ten times more trajectories than the original data, we only employed the tenth generated sample for each trajectory data point. We used the same number of Gaussian components for the implementation of the GMM as in ATRADA. Furthermore, to tackle the issue of noisy covariance matrices caused by high dimensionality, we utilized the singular value decomposition technique described in the original paper (Barratt et al., 2018).
@@ 4.4 Quantitative Evaluation
We performed discriminative and predictive tasks to quantitatively evaluate the quality of the generated aircraft trajectory data. For the discriminative task, we first trained a Transformer-based classifier to distinguish between actual and generated trajectories, labeling the former as real and the latter as fake. For the classification, we append a classification (CLS) token to the input sequence, an approach borrowed from Bidirectional Encoder Representations from Transformers (BERT) and Vision Transformer (ViT) (Devlin et al., 2019; Dosovitskiy et al., 2021). This CLS token is employed for classifying a given trajectory as either real or fake. We then computed the discriminative score $\left|\frac{1}{2}-\text{accuracy}\right|$ (DS-Classifier). A score closer to zero indicates that the classifier is unable to distinguish between actual and generated trajectories and performs no better than random guessing.
Additionally, we conducted an ATCo Turing test. In the ATCo Turing test, we engaged four professional ATCo who specialize in the airspace of ICN. They were asked to distinguish between real and generated trajectories. Each controller was shown ten trajectory snapshots. Each snapshot displayed both the horizontal view and the vertical profile of the trajectory. Half of these snapshots portrayed real trajectories, while the remaining five displayed generated trajectories. The discriminative score for the ATCo Turing test (DS-ATCo) was computed similarly to the classifier’s discriminative score.
Next, we trained a Transformer-based predictor on the generated aircraft trajectory dataset and evaluated its performance on an actual aircraft trajectory dataset, known as the Train on Synthetic, Test on Real (TSTR) approach. The predictor was designed to predict a future trajectory sequence over a time horizon of 120 seconds given a past trajectory sequence over a time horizon of 120 seconds. For training the predictor, we created pairs of past and future trajectory sequences, each spanning 20 points (120 seconds), by sliding a window with a size of one time step from the generated trajectories. We used the Mean Absolute Error (MAE) to assess the prediction accuracy (PS). Trajectory prediction is one of the major downstream applications of trajectory dataset augmentation, so it is important to validate the usefulness of the generated trajectories based on prediction accuracy.
@@ 4.5 Qualitative Evaluation
We applied t-SNE (Van der Maaten & Hinton, 2008) to the datasets of actual and generated aircraft trajectories to visualize how well they match each other. For a deeper analysis, we also applied it to the first and second derivatives of aircraft trajectories (i.e., velocity and acceleration), which allowed us to assess whether the generated trajectories adequately capture the dynamic behavior of aircraft over time. Additionally, we visualized the top view (latitude and longitude) of the generated trajectories, which allowed us to evaluate how well the generated trajectories satisfy various operational constraints, including the designated entry points into the airspace and the FAF that aircraft must pass through to prepare for landing.
@@ 5.1 Quantitative Evaluation
Table 2 presents the experimental results of the quantitative evaluation. The relative improvements (RIs) were calculated as $\left|s_{2}-s_{1}\right|/s_{2}$ to compare the best score $s_{1}$ to the second-best score $s_{2}$. ATRADA performed better than the other baselines and achieved higher RIs of 43.4%, 25.0%, and 36.0% across all three evaluation metrics. This indicates that ATRADA generates trajectories that are more indistinguishable from actual trajectories and more effectively retains their predictive characteristics.
TimeVAE and GMM also performed well. Specifically, the trajectories generated by TimeVAE achieve the second-best score for DS-Classifier metric and the third-best score for DS-ATCo metric. However, they tended to be less effective for predictive tasks compared to ATRADA and GMM. On the other hand, GMM generates trajectories that better preserve the predictive characteristics of the actual trajectories (i.e., lower PS) than TimeVAE, but its discriminative scores are not as strong as those of ATRADA and TimeVAE. SMOTE-I generates trajectories by interpolating two latent vectors and produced trajectories that are similar to actual trajectories, as indicated by strong scores for the DS-Classifier and DS-ATCo metrics. However, these generated trajectories were less useful for predictive tasks. Conversely, the opposite trend occurred with SMOTE-E, which extrapolates two latent vectors to generate synthetic trajectories.
Table 2: Experimental results for three quantitative evaluations. Bold and underlined values indicate the best and second-best performance, respectively.
Model DS-Classifier DS-ATCo PS
@@
ATRADA (Ours) 0.013 0.075 0.016
TimeGAN 0.414 0.350 0.084
TimeVAE 0.023 0.125 0.028
SMOTE-I 0.027 0.100 0.056
SMOTE-E 0.176 0.375 0.031
GMM 0.057 0.300 0.025
TCN-VAE 0.036 0.125 0.037
Original - - 0.009
RI 43.4% 25.0% 36.0%
@@ 5.2 Qualitative Evaluation
Figure 5 presents the t-SNE visualizations for the position ($1^{\text{st}}$ row), velocity ($2^{\text{nd}}$ row), acceleration ($3^{\text{rd}}$ row) for the generated trajectories from ATRADA, TimeVAE, and GMM, which produced good performance in the quantitative evaluation. The t-SNE visualizations for other baselines can be found in the Appendix A. The trajectories generated by ATRADA exhibit strong overlap with the actual trajectories across all three dimensions. In contrast, a significant portion of the trajectories from TimeVAE and GMM fall out of the distribution for acceleration, which indicates that they do not adequately capture the dynamic behavior of aircraft. Notably, TimeVAE also generates some out-of-distribution samples for position and velocity, which correspond to synthetic trajectories with arbitrary entry points into the airspace, as shown in Figure 6(c).
@@
Figure 6 presents top views of actual and generated trajectories, with entry points and the airport marked by black triangles and squares, respectively. The FAFs for different runway directions are also shown by red “+” symbols. ATRADA and TimeVAE generate smooth aircraft trajectories in the top view, while the trajectories generated by GMM exhibit highly irregular zig-zag patterns that are unrealistic for aircraft movements, as highlighted by the red boxes in Figure 6(d). Furthermore, some of the trajectories generated by TimeVAE spread across the airspace and fail to pass through the designated entry points or the FAF, whereas the trajectories from ATRADA consistently and precisely pass through these points.
@@ 6 Discussion
TimeGAN did not perform well in the task of aircraft trajectory generation despite being considered a state-of-the-art method for time-series data generation. We conjecture that this is due to its implicit likelihood training. Specifically, the generator of TimeGAN is designed to deceive the discriminator rather than accurately learn the probability distribution of the dataset. As a result, the trajectories generated by TimeGAN may fail to satisfy critical spatial constraints, such as designated entry points or STARs in the airspace.
In contrast, models with explicit likelihood training (ATRADA, TimeVAE, TCN-VAE, and GMM) performed better in generating trajectories that align with the spatial distribution of actual trajectories. However, TimeVAE and TCN-VAE still showed limited capability in accurately capturing the distribution of actual trajectories because they learn the dataset’s distribution indirectly through the reconstruction and regularization objectives. Conversely, ATRADA directly estimates the parameters of Gaussian distributions from the actual dataset, which potentially leads to more consistent and reliable trajectory generation.
The idea of using the context vectors in the latent space, rather than the original variables, has proven to offer statistical advantages in many prior works (Bengio et al., 2012; DeVries & Taylor, 2017). The results of this work are also consistent with these findings. Working with GMM in the latent space, rather than in the original input space, allows us to generate synthetic trajectories that are not only more similar to real trajectories but also more beneficial for the downstream tasks, such as trajectory prediction.
@@ 7 Conclusion
In this work, we have proposed a novel framework for aircraft trajectory dataset augmentation that leverages the representational capabilities of the Transformer as a sequence autoencoder and the probabilistic nature of the GMM as a generator. The effectiveness of ATRADA was validated using actual air traffic surveillance data. The results demonstrate that it generates more realistic aircraft trajectories that respect the operational constraints of real-world airspace. Given the highly constrained nature of air traffic operations due to safety concerns, the ability of ATRADA to generate trajectories that adhere to both aerodynamic and operational constraints is particularly advantageous. In summary, ATRADA not only exhibits better performance in generating realistic aircraft trajectories, but it could also be beneficial for various downstream tasks such as trajectory prediction and air traffic simulation.

### v3

@@ 4.2 Model configurations
We set the number of encoder layers to $L=3$, the hidden dimension to $D=128$, the number of heads to $h=4$, the feedforward dimension to $D_{ff}=512$, and the dropout probability to $p=0.2$. For training, we used the ADAM optimizer [23] with an initial learning rate of $10^{-4}$. The number of principal components $P$ for PCA was set to 22, and the number of Gaussian components $K$ for GMM was set to 32.
To determine $P$, we tested different values from 1 to 50 and selected 22 as the minimum number of components required to explain 99% of the cumulative variance. The scree plot for the PCA is presented in Figure 5. For the GMM, we evaluated the Bayesian information criterion (BIC) [24] for values of $K$ ranging from 1 to 50, and the optimal value was chosen as $K=32$, which achieved the lowest BIC, thus balancing the model likelihood and complexity. The BIC plot for different values of $K$ is shown in Figure 6.
Figure 5: Sensitivity analysis of cumulative explained variance with respect to the number of principal components in PCA.
Figure 6: Sensitivity analysis of BIC with respect to the number of Gaussian components of GMM.
@@ 4.3 Comparison baselines
We compared the proposed framework with several competitive baselines: TimeGAN [13], TimeVAE [14], SMOTE-Interpolation (SMOTE-I), SMOTE-Extrapolation (SMOTE-E) [25], GMM [3] and TCN-VAE [19]. We used the following publicly available source code for the implementation of the baselines (TimeGAN, TimeVAE, TCN-VAE and GMM). The source code for ATRADA will also be made publicly available.
For TimeGAN, the hidden dimensions of its three gated recurrent unit (GRU) layers were set to four times the size of the input features, which is consistent with the implementation described in the original TimeGAN paper [13]. The hidden dimensions of TimeVAE and TCN-VAE were configured to match the hidden dimension used in ATRADA. For both SMOTE-I and SMOTE-E [25], we set the number of nearest neighbors to 10, and the degrees of interpolation and extrapolation were set as 0.5. Given that SMOTE-I and SMOTE-E were tailored to generate 10 times more trajectories than the original data, we only used every tenth generated sample for each trajectory data point. We used the same number of Gaussian components for the implementation of the GMM as in ATRADA. Furthermore, to address the issue of noisy covariance matrices caused by high dimensionality, we used the singular value decomposition technique described in the original paper [3].
@@ 4.4 Quantitative evaluation
We performed discriminative and predictive tasks to quantitatively evaluate the quality of the generated aircraft trajectory data. For the discriminative task, we trained a Transformer-based classifier to distinguish between actual and generated trajectories. For the classification, we appended a classification (CLS) token to the input sequence, which is the approach used for bidirectional encoder representations from Transformers (BERT) and Vision Transformer (ViT) [26, 27]. This CLS token is used for classifying a given trajectory as either real or generated. We then computed the discriminative score as $\left|\frac{1}{2}-\text{accuracy}\right|$ (DS-Classifier). A score closer to zero indicates that the classifier is unable to distinguish between actual and generated trajectories and performs no better than random guessing.
We also conducted an ATC Turing test by engaging four professional ATCs who specialize in the airspace of ICN. They were asked to distinguish between real and generated trajectories among 10 trajectory snapshots, which displayed both a horizontal view and a vertical profile of the trajectory. Half of these snapshots portrayed real trajectories, while the remaining five displayed generated trajectories. The discriminative score for the ATC Turing test (DS-ATC) was computed similarly to DS-Classifier.
Next, we trained a Transformer-based predictor on the generated aircraft trajectory dataset and evaluated its performance on an actual aircraft trajectory dataset, which is known as the “train on synthetic, test on real” (TSTR) approach. The predictor was designed to predict a future trajectory sequence over a time horizon of 120 seconds given a past trajectory sequence over a time horizon of 120 seconds. For training the predictor, we created pairs of past and future trajectory sequences spanning 20 points each (120 seconds) by sliding a window with a size of one time step from the generated trajectories. We used the mean absolute error (MAE) to assess the prediction accuracy (PS). Trajectory prediction is one of the major downstream applications of trajectory dataset augmentation, so it is important to validate the usefulness of the generated trajectories based on prediction accuracy.
We also directly evaluated the quality of the generated trajectories and assessed their statistical similarity to the actual data by calculating the Kullback–Leibler divergence (KLD) between the actual and synthetic trajectory distributions. KLD quantifies the discrepancy between the probability distribution of the actual data and that of the generated trajectories and is defined as follows:
$\text{KLD}(P\parallel Q)=\sum_{i}P(x_{i})\log\frac{P(x_{i})}{Q(x_{i})}$ (4)
where $P$ and $Q$ denote the probability distributions of the actual and generated trajectories, respectively.
@@ 4.5 Qualitative evaluation
We applied t-distributed stochastic neighbor embedding (t-SNE) [28] to the datasets of actual and generated aircraft trajectories to visualize how well they match each other. For deeper analysis, we also applied it to the first and second derivatives of aircraft trajectories (velocity and acceleration) to assess whether the generated trajectories adequately capture the dynamic behavior of aircraft over time. We visualized the top view (latitude and longitude) of the generated trajectories to evaluate how well the generated trajectories satisfy various operational constraints, including the designated entry points to the airspace and the FAF that aircraft must pass through to prepare for landing.
@@ 5.1 Quantitative evaluation
Table 2 presents the experimental results of the quantitative evaluation. Each column shows the scores of ATRADA and the baseline models for each evaluation metric. DS-Classifier, DS-ATC, and KLD require comparison between actual and generated trajectory data, so reporting these scores for the original dataset is not applicable, and they were omitted from the table.
The relative improvements (RIs) of ATRADA over the second-best baseline models are provided at the bottom of the table. RI was calculated as $\left|s_{2}-s_{1}\right|/s_{2}$ , where $s_{1}$ denotes the best score achieved by the proposed method (indicating the lowest error), and $s_{2}$ is the second-best score among the baselines. In other words, the RI represents the relative error reduction of ATRADA compared to the second-best baseline for each evaluation metric.
ATRADA outperformed all baseline models and achieved RIs of 43.4%, 25.0%, 36.0%, and 17.1% across the four evaluation metrics compared to TimeVAE, SMOTE-I, GMM, and GMM, respectively. This result indicates that ATRADA generates trajectories that are more indistinguishable from actual trajectories and more effectively preserves their predictive characteristics. Furthermore, the KLD results confirm that ATRADA generates trajectories that are statistically closer to the actual trajectories.
TimeVAE and GMM also performed well. Specifically, the trajectories generated by TimeVAE achieve the second-best score for DS-Classifier metric and the third-best score for DS-ATC metric. However, they tended to be less effective for predictive tasks compared to ATRADA and GMM. On the other hand, GMM generated trajectories that better preserve the predictive characteristics of the actual trajectories (lower PS) than TimeVAE, but its discriminative scores were not as strong as those of ATRADA and TimeVAE. SMOTE-I generates trajectories by interpolating two latent vectors and produced trajectories that are similar to actual trajectories, as indicated by strong scores for the DS-Classifier and DS-ATC metrics. However, these generated trajectories were less useful for predictive tasks. Conversely, the opposite trend occurred with SMOTE-E, which extrapolates two latent vectors to generate synthetic trajectories.
Table 2: Experimental results for four quantitative evaluations. Bold and underlined values indicate the best and second-best performance, respectively.
Model DS-Classifier DS-ATC PS KLD
@@
ATRADA (Ours) 0.013 0.075 0.016 44.001
TimeGAN 0.414 0.350 0.084 2706.021
TimeVAE 0.023 0.125 0.028 75.472
SMOTE-I 0.027 0.100 0.056 60.525
SMOTE-E 0.176 0.375 0.031 762.927
GMM 0.057 0.300 0.025 53.061
TCN-VAE 0.036 0.125 0.037 673.747
Original - - 0.009 -
RI 43.4% 25.0% 36.0% 17.1%
@@ 5.2 Qualitative evaluation
Figure 7 presents the t-SNE visualizations for the position ($1^{\text{st}}$ row), velocity ($2^{\text{nd}}$ row), acceleration ($3^{\text{rd}}$ row) for the generated trajectories by ATRADA, TimeVAE, and GMM, which produced good performance in the quantitative evaluation. The t-SNE visualizations for other baselines can be found in Appendix A. The trajectories generated by ATRADA exhibit strong overlap with the actual trajectories across all three dimensions. In contrast, a significant portion of the trajectories from TimeVAE and GMM fall out of the distribution for acceleration, which indicates that they do not adequately capture the dynamic behavior of aircraft.
Notably, TimeVAE also generated some out-of-distribution samples for position and velocity, which correspond to synthetic trajectories with arbitrary entry points into the airspace, as shown in Figure 8(c). Note that during the process of projecting each trajectory into a 2-demensional space using t-SNE, the x- and y-axes are determined by the distribution of the combined dataset (i.e., the actual trajectories together with the synthetic trajectories generated by each method). As a result, although the same set of actual trajectories is used across all three methods (ATRADA, TimeVAE, and GMM), their visual shapes appear different in each subfigure of Figure 7.
Figure 8 presents top views of actual and generated trajectories, with entry points and the airport marked by black triangles and squares, respectively. The FAFs for different runway directions are also shown by red symbols (“+”). ATRADA and TimeVAE both generated smooth aircraft trajectories in the top view. However, TimeVAE generated trajectories that spread widely across the airspace and often failed to pass through the designated airspace entry points or the FAFs, as highlighted by the red boxes in Figure 8(c). Trajectories that do not pass through the FAFs may imply landings between runways or outside valid runway boundaries, which cannot occur in actual operations.
Similarly, GMM generates trajectories that often fail to pass through the FAFs and exhibit highly irregular zig-zag patterns that are unrealistic for aircraft movements, as highlighted by the red boxes in Figure 8(d). In contrast, ATRADA consistently generated trajectories that precisely pass through these points, which demonstrates their realism and their potential utility for downstream tasks.
@@ 6 Discussion
TimeGAN did not perform well in the task of aircraft trajectory generation despite being considered a state-of-the-art method for time-series data generation. We conjecture that this resulted from its implicit likelihood training. Specifically, the generator of TimeGAN is designed to deceive the discriminator rather than accurately learn the probability distribution of the dataset, so the trajectories generated may fail to satisfy critical spatial constraints, such as designated entry points or STARs in the airspace.
In contrast, the models with explicit likelihood training (ATRADA, TimeVAE, TCN-VAE, and GMM) performed better in generating trajectories that align with the spatial distribution of actual trajectories. However, TimeVAE and TCN-VAE still showed limited capability in accurately capturing the distribution of actual trajectories because they learn the dataset’s distribution indirectly through reconstruction and regularization objectives. Conversely, ATRADA directly estimates the parameters of Gaussian distributions from the actual dataset, which can lead to more consistent and reliable trajectory generation.
The idea of using the context vectors in the latent space rather than the original variables has shown statistical advantages in many prior works [29, 25]. The results of this work are also consistent with these findings. Working with GMM in the latent space rather than in the original input space allows us to generate synthetic trajectories that are not only more similar to real trajectories but also more beneficial for the downstream tasks, such as trajectory prediction.
@@ 7 Conclusion
We have proposed a novel framework for aircraft trajectory dataset augmentation that leverages the representational capabilities of the Transformer as a sequence autoencoder and the probabilistic nature of the GMM as a generator. The effectiveness of ATRADA was validated using actual air traffic surveillance data. The results demonstrated that it generates more realistic aircraft trajectories that respect the operational constraints of real-world airspace. Given the highly constrained nature of air traffic operations due to safety concerns, the ability of ATRADA to generate trajectories that adhere to both aerodynamic and operational constraints is particularly advantageous. In summary, ATRADA exhibits better performance in generating realistic aircraft trajectories, and could be beneficial for various downstream tasks, such as trajectory prediction and air traffic simulation.
@@ Appendix A Additional Visualizations
Conflict-of-Interest Statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.